OpenStudy (anonymous):

Simplify each expression:? 1) 8(-b-4) 2) (1-7n)x5 3) -6(x-4) 4) (1-7n)x5

OpenStudy (nubeer):

1. 8(-b-4) = -8b -32

OpenStudy (anonymous):

open the brackets by multiple each term with the number outside the bracket

OpenStudy (nubeer):

same way u have to do the rest just multiply the number outside with the numbers inside.

OpenStudy (anonymous):

1) 8(-b-4) = -8b - 32 2) (1-7n)x5 = 5 - 35n 3) -6(x-4) = -6x + 24 4) (1-7n)x5 = 5 - 35n

OpenStudy (anonymous):

1. -8b +32 2. 5 - 35n 3. -6x + 24 4. 5 - 35n

OpenStudy (anonymous):

are the simplifications.

OpenStudy (anonymous):

just take care of -ve sign